What is the shortest wavelength of the continuum p
Answer #1

we have than the energy for a wavelength is E = \frac{hc}{\lambda } h = planck ctte = 4.18x10-18KeV/s c= 2.9979x108m/s and the we have E = 1.2397x10-9/\lambda =) \lambda  =1.2397x10-9/E =)\lambda  = 1.2397x10-9Kev m/97Kev =) \lambda  = 1.27x10-11m

\lambda  = 1.27x10-11m